# Image-Based Virtual Try-On: A Survey

### v2

@@ 3.2 Evaluation Criteria
IS: Inception Score [137] is a metric based on Inception network, which is originally used to evaluate generation models in terms of clarity and diversity. The calculation of IS only involves generated images, and the prediction probability of generated images is used to describe the performance. Clarity refers to more certain probability of belonging to one class, e.g., the prediction entropy is smaller. Diversity involves a set of generated images, where the predictions are evenly distributed across all classes indicate better diversity. The higher value of IS indicate the better generation performance. However, different from the original generation tasks, the generation target for virtual try-on is a fixed class such as human photos. It is not fair enough for this task, and FID becomes an alternative way which compares the distributions of generated images and ground truth.
@@ 4 Experimental Results and Analysis
Due to the lack of unified evaluation and comprehensive analysis, we run existing open-source codes in image-based virtual try-on area and evaluate them with the same referees. In this section, experimental results are given as quantitative results, qualitative results and user study. Besides the observations towards these results, suggestions in terms of method design are also analyzed.
@@ 4.1 Quantitative Results
We consider two testing conditions to quantitatively evaluate current representative approaches, i.e., paired condition and unpaired condition. Ground truth exists for paired condition while for unpaired condition the clothes in clothing image and person image are different. Unpaired condition shows a practical scenario but this situation lacks ground truth. With the human parsing method SCHP [142], partial comparisons are proposed for the first time, which separately evaluate the performance of clothing warping (denoted as “clothing”) and generation of non-try-on area (denoted as “person”). More details can be found in the supplement.
The quantitative results of representative methods in the order of publication year are shown in Fig. 7. As unpaired condition lacks ground truth, it is hard to evaluate the generation quality of full body for SSIM, LPIPS and Semantic Score. From the results, we have the following observations:
  • Except the results of “Paired Clothing” in terms of SSIM and IS, the overall performance gets better along the development. The poor SSIM performance of clothing warping can be attributed to the damage of clothing patterns.
@@
(a) SSIM $\uparrow$
(b) IS $\uparrow$
(c) FID $\downarrow$
(d) LPIPS $\downarrow$
(e) Semantic Score $\downarrow$
Figure 7: Quantitative results. The vertical axis indicates the testing condition is paired or unpaired and the evaluation target is clothing, non-try-on area (denoted as “person”) or all together. For method symbol, the cross symbol indicates the clothing warping strategy is flow while the rectangle symbol represents TPS-based method. *Both HR-VTON and VITON-HD are tested using high-resolution ($1024\times 768$) datasets. To compare with other low-resolution models, we down-sample the high-resolution results to low resolution ($256\times 192$). To illustrate the impact of image resolution on evaluation metrics, we denote the scores of high-resolution results with gray points.
@@ 4.2 Qualitative Results
The qualitative results are displayed in Fig. 8 by showing several cases of image-based virtual try-on. The $1^{st}$ row shows an easy task where the in-shop clothing has the same type as the source image and only contains plain color. Previous methods perform comparatively and the results are satisfying. When changing a piece of clothes with sleeves to a non-sleeve type (e.g., the $2^{nd}$ and $3^{rd}$ row), the sleeves are usually completely removed. However, as shown in the $4^{th}$ row, it is difficult to perfectly fit the sleeves to the person arms. When focusing on the person neck of try-on images, the clothing style is affected by the performance of human parser. The preservation of clothing texture such as special patterns (e.g., the $3^{rd}$ and $5^{th}$ row) and stripes (e.g., the $6^{th}$ row) could be further improved. Person shape (e.g., $3^{rd}$ row) has little effects while challenging pose (e.g., the last row) will increase the difficulty of try-on. Challenging cases will be further discussed in Sec. 5.
Figure 8: Qualitative results. We show the visual results with variations of person poses, shapes and clothing patterns. Please refer to Sec. 4.2 for detailed analysis.
From qualitative results, Style-Flow-VTON [38] performs best under low resolution. For earlier works such as VITON [21], SP-VTON [25] and CP-VTON [23] models, the focus is on the processing of the try-on part. When trying on simple clothes, the center area of the clothes has a better generation effect, but the clothes and person parts in the non-try-on area are blurred. LM-VTON [24] and CP-VTON+ [22] have improved the above problems to some extent by preserving the image contents in the non-try-on area, but also contain artifacts. Although DCTON [71] adopts a different model framework from the VITON series, its performance is close to that of LM-VTON and CP-VTON+. The quality of the results generated by PFAFN [37], Style-Flow-VTON [38] and SDAFN [74] gets better, where the boundary between the body area and the clothing area is more natural and clear. It mainly attributes to the warping ability of flow. VITON-HD [39] and HR-VITON [31] are two methods designed for high-resolution virtual try-on, and the results of HR-VITON contain more details.
@@ 4.4 Discussion on Method Design
  • Pipeline perspective. The structures of existing pipelines are classified as shown in Fig. 2, and the specific pipeline adopted by each approach is illustrated in Table I. It could be found that there is no obvious pipeline preference in the development of image-based virtual try-on methods. From both the quantitative and the qualitative results, the dual-path pipeline (i.e., type \@slowromancapvii@) achieves superior performance where the dual paths facilitate each other to optimize the generation performance.
  • Clothing warping perspective. Existing warping approaches adopted by representative methods are shown in Table I. As illustrated in Sec. 2.5, spatial transformation network (STN) usually cooperates well with thin plate spline (TPS) transformation, and TPS plays a dominating role in the development of clothing warping methods. The performance of clothing warping approaches is reflected as the results of “paired clothing” in Fig. 7, where the rectangle symbol represents TPS-based method and the cross mark denotes flow-based method. Comparatively, the warping method of flow estimation shows superior performance with more flexible transformation.
  • Try-on perspective. The final try-on performance depends on all procedures including try-on indication, cloth warping and try-on. The generation quality gets better with the development of generative model such as StyleGAN [40, 41, 42] and Diffusion Model [55, 143].
Figure 10: Performance under challenging poses. $1^{st}$ row: the results in a classic arms-crossed posture. $2^{nd}$ row: a more challenging case in arms occlusion where the color of sleeves is the same as torso. $3^{rd}$ row: putting on a piece of long-sleeved clothes under the arms-crossed posture. $4^{th}$ row: the fine-grained details of bent wrist cannot be well recovered.
@@ 5 Unresolved Issues
From previous qualitative results, we can find that try-on results with clear and fine-grained details remain challenging. Apart from this, in this section we further show examples for issues caused by human parsing. Human parsing plays a vital role in image-based virtual try-on, but also causes the main issues due to imperfect parsing ability.
@@ 5.1 Challenging Poses
Figure 12: The generation of body with challenging poses is alleviated with the help of large-scale models. The try-on results are generated using a slightly modified PBE [55] model as the baseline for validating the virtual try on task based on large models.
For problem a, from Fig. 10 we can find that most methods can tell the front or back position correctly (e.g., the results in $1^{st}$ row) except the ambiguous clothing with plain color (e.g., the results of PFAFN [37] and SDAFN [74] in the $2^{nd}$ row). For problem b, the crossed arms are not well solved. HR-VTON[31] and VITON-HD[39] mitigate this problem by optimizing the misaligned parts of the human body and clothing within the generator. The results in the $3^{rd}$ row reflect the model’s ability of handling problem c, where the character aims to wear a piece of long-sleeved clothes. Most methods fail to correctly put on sleeves due to the limitation of clothing warping and try-on module. In addition, the arms-crossed pose contains the crossing of the fingers and the arm, which requires high granularity generation ability of the model. Take the $4^{th}$ row for example, current representation such as Openpose [60], Densepose [58] and semantic segmentation map [56, 142] cannot label the fine-grained pose at the wrist (Fig. 11), which fundamentally leads to the poor generation quality.
With the advent of large-scale image generation models [55, 50, 144], the issue of blurred body has been significantly improved. Fig. 12 shows the results that we utilize PBE [55] to inpaint the target clothes to human body. The postures are correctly restored. Thanks to the powerful zero-shot generation capabilities of large-scale models, simple posture features are sufficient to guide the model to generate the correct body. However, it will be pointed out in Sec. 6 that current large-scale models are not fully controllable in terms of the generated contents.
Figure 13: Changed clothing styles. The styles of target clothes cannot be preserved due to introducing original clothing information in person representation or clothing bias in the training dataset.
@@ 5.2 Limited Human Parsing
Clothing style is an important aspect to show the performance of virtual try-on, but the current models has poor performance on unconventional styles. From top to bottom in Fig. 13 are short T-shirts, loose-try-on tops, long T-shirts and light tops. Compared with the correct wearing state of the corresponding clothing in the rightmost column of each row, the try-on results generated by the model have obvious biases, and the model tends to make the clothing fit perfectly on the upper body of the person body, presenting a slimming effect. This is unfair for clothing with loose styles or special designs in length, such as the results in the first and third rows, where the short T-shirts are forcibly stretched to the normal length of T-shirts, and the long T-shirts are forcibly reduced/truncated to the normal length of T-shirts. Essentially, the model does not really understand the clothing to be tried on, and it only deforms the clothing image and fills it in the original clothing area for replacement. The main reason attributes to that the model is trained with paired data where target clothes and the clothes on person are actually the same one, and in reality it is difficult to collect unpaired data. Clothing-agnostic person representation still deserve exploration.
Figure 14: Artifacts caused by noisy semantic parsing.
@@ 5.3 Limited Clothing Parsing
It is worth mention that for large-scale image generation models, e.g., our implemented PBE for virtual try-on, only the character image and clothing image are needed as conditions, and a mask is used to guide the area to be tried on. Due to the strong feature parsing ability and scalability of large image generation models, the dependency on parsing will be also alleviated.
@@ 6 Future Work
Clothing-agnostic person representation. As the training data is paired where the person already wears target clothes, it will harm the generalization ability if person representation contains clothes information. Current representations $\mathcal{P}_{\begin{subarray}{c}3,6,7,10,11\end{subarray}}$ in Fig. 4 contains clues from the original clothes such as clothing contour, which will affect the clothing style of target clothes. Totally deleting the body area (i.e., $\mathcal{P}_{\begin{subarray}{c}4,12\end{subarray}}$ will get rid of original clothes but lose person characteristics such as skin color. Constructing triplets where a person wears different clothes could be a solution, but current methods [36, 37, 38] still rely on parser-based methods to construct the triplet, which inherits the limitation of human parser. Future direction could try to construct training triplets with parser-free approaches such as recent large-scale models, or totally delete original clothes but keep a sampler at exposed skin.
@@
Controllable generation with large-scale model. Large-scale image generation models show powerful generalization and high quality, which is appealing for virtual try-on tasks. However, current models only meet semantic-level control such as satisfying the description from text. More precise and strict control has not been solved such as preserving the original patterns (e.g., Fig 16). To enhance the adaptability of large models for virtual try on tasks, we utilize ControlNet [54] to fine-tune the pre-trained large model [55] with target clothing image as condition. As shown in Fig. 16, the results are natural and realistic but the original patterns are not preserved. In the future, training a specific large-scale model for virtual try-on and preserve original clothing patterns deserve investigation.
Figure 16: Generated try-On images by finetuning the stable diffusion model with ControlNet [54]. The generated results are realistic but the clothing patterns are not well preserved.
@@
Specialized evaluation metrics. Existing evaluation metrics are not specialized for virtual try-on tasks, and do not reach a consensus with user study. In this paper, a new criteria is proposed with advanced pre-trained model and specifically evaluate the semantic information for try-on and non-try-on areas. Future attention could be paid to assessing the clarity and realism of clothing patterns and the preservation of person characteristics.
@@
Scenario-specific try-on. Real-world try-on is limited to dressing rooms, while try-on effects differ in different scenarios such as indoor or outdoor, different seasons and different occasions. Virtual try-on has advantages in synthesizing the background, which provides imaginary space for users to consider the suitability of target clothes. Furthermore, generating videos to showcase the effects in different scenarios would be more engaging.
Appendix A Implementation Details
@@
The code link for comparison methods and the implementation of evaluation metrics will be given at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.
Appendix B Evaluation Criteria
Previous methods are quantitatively evaluated in terms of testing conditions and areas. Fig. 17 shows two testing conditions, i.e., paired and unpaired. Paired situation refers to that we have a pair of clothing image and image of person wearing the same clothes. Such data are abundant on the Internet and usually used for training by masking the person image to obtain clothing agnostic person representation. Unpaired situation is close to practical scenario where the clothing in the person image is different from target clothing, where we do not have the ground truth for try-on image. Due to the lack of ground truth, the evaluation criteria such as SSIM, LPIPS and semantic score that rely on paired data do not have unpaired results.
Figure 17: Two testing conditions are involved for evaluation. Paired evaluation refers to the condition that the clothing image and person image are paired and the clothing-agnostic image is parsed from the person image. Unpaired evaluation refers to that the clothes in the clothing image and person image are different, so there is no ground truth for the try-on result.
We for the first time separately evaluate the generation quality of different areas, i.e., the warped clothing area, the non-try-on area, and the whole area, which are respectively denoted as “Clothing”, “Person” and “All”. Fig. 18 shows several examples for different evaluation areas.
Figure 18: Three testing conditions (“All”, “Person” and “Clothing”). With the help of a parser [142], we separate the generated image to non-try-on area and clothing area. The deleted areas are replaced with black pixel.
Appendix C Quantitative Results
Fig. 7 draws the quantitative results in charts to better show the relative performance between methods. The original numerical results are shown in Tab. III, IV, V, VI and VII.
Appendix D More Visual Results
All of the testing results will be given at https://github.com/little-misfit/Survey-Of-Virtual-Try-On. In terms of clothing style and texture, we show more visual results in Fig. 19.
TABLE III: SSIM $\uparrow$
Model Paired Clothing Paired Person Paired All Unpaired Person
VITON[21] 0.660 0.848 0.824 0.800
CP-VTON [23] 0.661 0.787 0.763 0.751
SP-VTON [25] 0.663 0.797 0.778 0.759
CP-VTON+[22] 0.659 0.842 0.814 0.784
LM-VTON [24] 0.662 0.886 0.869 0.864
PFAFN [37] 0.660 0.920 0.920 0.860
DCTON[71] 0.657 0.887 0.873 0.840
VITON-HD[39] 0.664 0.860 0.824 0.829
Style-Flow-VITON[38] 0.659 0.921 0.924 0.868
SDAFN[74] 0.661 0.890 0.887 0.832
HR-VTON[31] 0.660 0.873 0.833 0.832
TABLE IV: LPIPS $\downarrow$
Model Paired Clothing Paired Person Paired All Unpaired Person
VITON[21] 0.166 0.135 0.155 0.193
CP-VTON[23] 0.184 0.158 0.179 0.211
SP-VTON[25] 0.207 0.185 0.205 0.229
CP-VTON+[22] 0.165 0.119 0.134 0.192
LM-VTON[24] 0.203 0.125 0.133 0.142
PFAFN[37] 0.137 0.090 0.076 0.174
DCTON[71] 0.171 0.112 0.111 0.177
VITON-HD[39] 0.183 0.125 0.154 0.173
Style-Flow-VITON[38] 0.141 0.090 0.073 0.166
SDAFN[74] 0.152 0.112 0.098 0.191
HR-VTON[31] 0.167 0.116 0.144 0.175
TABLE V: Semantic Score $\downarrow$
Model Paired Clothing Paired Person Paired All Unpaired Person
VITON[21] 5.773 4.496 4.461 5.086
CP-VTON[23] 6.535 6.457 5.730 6.613
SP-VTON[25] 6.176 5.559 5.000 5.441
CP-VTON+[22] 6.121 5.672 5.211 6.430
LM-VTON[24] 7.125 4.957 4.934 4.891
PFAFN[37] 5.051 4.176 3.695 4.773
DCTON[71] 6.293 4.707 4.707 5.387
VITON-HD[39] 5.426 2.998 3.883 3.629
Style-Flow-VITON[38] 5.113 4.043 3.775 4.574
SDAFN[74] 5.266 4.469 4.023 5.012
HR-VTON[31] 4.656 2.951 3.549 3.652
TABLE VI: FID $\downarrow$
Model Paired Clothing Paired Person Paired All Unpaired Clothing Unpaired Person Unpaired All
VITON[21] 72.289 30.192 30.144 86.379 32.155 31.236
CP-VTON[23] 72.856 26.963 24.969 68.908 26.827 27.048
SP-VTON[25] 82.710 36.641 37.855 76.419 32.107 29.927
CP-VTON+[22] 64.684 19.512 18.688 75.407 21.290 23.450
LM-VTON[24] 90.996 23.588 20.417 64.762 17.800 17.051
PFAFN[37] 45.065 15.607 8.608 49.649 16.701 11.431
DCTON[71] 71.489 18.452 16.285 79.167 19.903 20.416
VITON-HD[39] 39.241 7.658 14.747 39.896 8.012 15.223
Style-Flow-VITON[38] 44.215 15.718 7.395 48.256 16.419 10.158
SDAFN[74] 45.835 16.056 9.707 49.902 16.467 12.226
HR-VTON[31] 33.847 10.459 13.878 35.070 11.058 14.984
TABLE VII: IS $\uparrow$
Model Paired Clothing Paired Person Paired All Unpaired Clothing Unpaired Person Unpaired All
VITON[21] 4.475 2.479 2.764 4.281 2.374 2.697
CP-VTON[23] 4.332 2.451 2.941 4.072 2.384 2.759
SP-VTON[25] 4.540 2.501 2.529 4.279 2.385 2.663
CP-VTON+ [22] 4.268 2.645 3.148 4.122 2.565 3.074
LM-VTON [24] 4.603 2.693 3.137 4.056 2.606 2.750
PFAFN[37] 4.144 2.679 2.978 4.197 2.632 2.894
DCTON[71] 4.390 2.628 2.822 4.344 2.608 2.743
VITON-HD[39] 3.958 2.624 2.979 4.141 2.589 2.962
Style-Flow-VITON[38] 4.150 2.676 3.100 4.133 2.609 2.991
SDAFN[74] 4.084 2.633 3.001 4.012 2.597 2.917
HR-VTON[31] 3.633 2.736 3.114 3.609 2.650 3.032
Figure 19: More virtual try-on results. The left-most column shows the changing of clothing style and texture. For simplicity, clothing style refers to short sleeve or long sleeve, and clothing texture contains plain color, pattern, stripe and lace.
Appendix E Demographics of Participants
For user study, we released a public link to collect questionnaires. Each questionnaire consists of 22 single-choice questions in regards of the satisfactory towards try-on results, where 2 random examples of 11 methods are selected. The participants voluntarily fill out the questionnaire and receive the bonus upon completing the survey. Demographics of participants are shown in Fig. 20, where the gender ratio is balanced, with a majority of them being young people who engage in online shopping.
(a) Gender
(b) Age
(c) Frequency
Figure 20: Demographics of participants in regards of gender, age and frequency of shopping online.

### v3

@@ 3.2 Evaluation Criteria
IS: Inception Score [105] is a metric based on Inception network, which is originally used to evaluate generation models in terms of clarity and diversity. The calculation of IS only involves generated images, and the prediction probability of generated images is used to describe the performance. Clarity refers to more certain probability of belonging to one class, e.g., the prediction entropy is smaller. Diversity involves a set of generated images, where the predictions are evenly distributed across all classes indicate better diversity. The higher value of IS indicate the better generation performance. However, different from the original generation tasks, the generation target for virtual try-on is a fixed class such as human photos. It is not fair enough for this task (therefore not adopted in our unified evaluation), and FID becomes an alternative way which compares the distributions of generated images and ground truth.
@@
Figure 7: Quantitative results on VITON. The vertical axis indicates the testing condition is paired or unpaired and the evaluation target is clothing, non-try-on area (denoted as “person”) or all together. For method symbol, the cross symbol indicates the clothing warping strategy is flow while the triangle symbol represents TPS-based method. *Both HR-VTON and VITON-HD are tested using high-resolution ($1024\times 768$) datasets. To compare with other low-resolution models, we down-sample the high-resolution results to low resolution ($256\times 192$). To illustrate the impact of image resolution on evaluation metrics, we denote the scores of high-resolution results with gray points.
@@ 4 Experimental Results and Analysis
Due to the lack of unified evaluation and comprehensive analysis, we run existing open-source codes in image-based virtual try-on area and evaluate them with the same referees. In this section, experimental results are given as quantitative results, qualitative results and user study. Besides the observations towards these results, suggestions in terms of method design are also analyzed. Previous methods adopt VITON while recent methods use VITON-HD for validation, so we keep the original dataset adopted by representative methods for fair comparison and use two methods (i.e., HR-VITON [61] and VITON-HD [18]) as a bridge evaluated by both datasets.
Figure 8: Quantitative results on VITON-HD.
@@ 4.1 Quantitative Results
We consider two testing conditions to quantitatively evaluate current representative approaches, i.e., paired condition and unpaired condition. Ground truth exists for paired condition while for unpaired condition the clothes in clothing image and person image are different. Unpaired condition shows a practical scenario but this situation lacks ground truth. With the human parsing method SCHP [67], partial comparisons are proposed for the first time, which separately evaluate the performance of clothing warping (denoted as “clothing”) and generation of non-try-on area (denoted as “person”). The quantitative results of representative methods in the order of publication year are shown in Fig. 7 and 8. As unpaired condition lacks ground truth, it is hard to evaluate the generation quality of full body for SSIM, LPIPS and Semantic Score. From the results, we have the following observations:
Figure 9: Qualitative results on VITON. We show the visual results with variations of person poses, shapes and clothing patterns.
@@
  • With VITON-HD and HR-VITON as a bridge, by comparing Fig. 7 and 8 we can find that the overall generation quality of Diffusion-based methods outperforms GAN-based methods.
@@ 4.2 Qualitative Results
The qualitative results are displayed in Fig. 9 and 10 by showing several cases of image-based virtual try-on. The $1^{st}$ row in both figures shows an easy task where these representative methods are good at, and we can find that recent methods (i.e., in Fig. 10) generate results with high-fidelity and fine-grained details. The preservation of clothing texture such as special patterns and stripes (e.g., the $2^{nd}$, $3^{rd}$ and $4^{th}$ rows) could be further improved. GAN-based methods usually generate blurry results while Diffusion-based methods face challenges in controllability. Person shape (e.g., $4^{th}$ row) has little effects while challenging pose (e.g., the last row) will increase the difficulty of try-on. Challenging cases will be further discussed in Sec. 5.
Figure 10: Qualitative results on VITON-HD. We show the visual results with variations of person poses, shapes and clothing patterns.
From qualitative results, Style-Flow-VTON [40] performs best under low resolution in Fig. 9. For earlier works such as VITON [37], SP-VTON [111] and CP-VTON [119] models, the focus is on the processing of the try-on part. When trying on simple clothes, the center area of the clothes has a better generation effect, but the clothes and person parts in the non-try-on area are blurred. LM-VTON [75] and CP-VTON+ [82] have improved the above problems to some extent by preserving the image contents in the non-try-on area, but also contain artifacts. Although DCTON [30] adopts a different model framework from the VITON series, its performance is close to that of LM-VTON and CP-VTON+. The quality of the results generated by PFAFN [31], Style-Flow-VTON [40] and SDAFN [5] gets better, where the boundary between the body area and the clothing area is more natural and clear. It mainly attributes to the warping ability of flow. VITON-HD [18] and HR-VITON [61] are two methods designed for high-resolution virtual try-on, and the results of HR-VITON contain more details.
Overall, the results of Diffusion-based methods (in Fig 10) show clear appearance. GP-VTON [126] is a GAN-based method, but shows competitive performance, which attributes to its clothing warping strategy of separately warping partitions. DCI-VTON [34] first warps clothing and then use diffusion model to refine the try-on results, which both controls the fidelity of original clothes and shows clear appearance. Although Diffusion-based methods (e.g., Ladi-VTON [84], CAT-DM [136] and Stable VITON [56]) show clear appearance, the controllability in clothing color and texture can still be improved.
@@ 4.4 Discussion on Method Design
  • Pipeline perspective. The structures of existing pipelines are classified as shown in Fig. 2, and the specific pipeline adopted by each approach is illustrated in Table 1. It could be found that there is no obvious pipeline preference in the development of image-based virtual try-on methods, and recent methods tend to design one-stage pipeline (i.e., type \@slowromancapi@) with large-scale models. Except recent diffusion-based methods, the dual-path pipeline (i.e., type \@slowromancapvii@) achieves superior performance where the dual paths facilitate each other to optimize the generation performance.
  • Clothing warping perspective. Existing warping approaches adopted by representative methods are shown in Table 1. As illustrated in Sec. 2.5, spatial transformation network (STN) usually cooperates well with thin plate spline (TPS) transformation, and TPS plays a dominating role in the development of clothing warping methods. The performance of clothing warping approaches is reflected as the results of “paired clothing” in Fig. 7, where the triangle symbol represents TPS-based method and the cross mark denotes flow-based method. Comparatively, the warping method of flow estimation shows superior performance with more flexible transformation. Additionally, Table 1 shows that most diffusion-based methods do not contain an explicit clothing warping module, indicating that this module becomes not necessary with the improved performance of generation network.
  • Try-on perspective. The final try-on performance depends on all procedures including try-on indication, cloth warping and try-on. The generation quality gets better with the development of generative model such as StyleGAN and Diffusion model.
Figure 13: Performance under challenging poses on VITON. $1^{st}$ row: the results in a classic arms-crossed posture. $2^{nd}$ row: a more challenging case in arms occlusion where the color of sleeves is the same as torso. $3^{rd}$ row: the fine-grained details of bent wrist cannot be well recovered. $4^{th}$ row: putting on a piece of long-sleeved clothes under the arms-crossed posture.
Figure 14: Performance under challenging poses on VITON-HD.
@@ 5 Unresolved Issues
As illustrated in Table 1, i.e., person representation, most methods rely on human parsing. Human parsing plays a vital role in image-based virtual try-on, but also causes the main issues due to imperfect parsing ability. In this section, we further show examples for issues caused by human parsing.
@@ 5.1 Challenging Poses
For problem a, from Fig. 13 we can find that most methods can tell the front or back position correctly (e.g., the results in $1^{st}$ row) except the ambiguous clothing with plain color (e.g., the results of PFAFN [31] and SDAFN [5] in the $2^{nd}$ row). For problem b, the crossed arms are not well solved. HR-VTON [61] and VITON-HD [18] mitigate this problem by optimizing the misaligned parts of the human body and clothing within the generator.
Figure 15: Changed clothing styles (VITON). The styles of target clothes cannot be preserved due to introducing original clothing information in person representation or clothing bias in the training dataset.
The results in the $4^{th}$ row reflect the model’s ability of handling problem c, where the character aims to wear a piece of long-sleeved clothes. Most methods fail to correctly put on sleeves due to the limitation of clothing warping and try-on module. In addition, the arms-crossed pose contains the crossing of the fingers and the arm, which requires high granularity generation ability of the model. Take the $3^{rd}$ row for example, current representation such as Openpose [13], Densepose [35] and semantic segmentation map [33, 67] cannot label the fine-grained pose at the wrist (Fig. 12), which fundamentally leads to the poor generation quality.
With the advent of large-scale image generation models [129, 142, 16], the issue of blurred body has been significantly improved. As shown by the results of Diffusion-based methods in Fig. 14, the position of arms and details of figures are well perceived. However, it should be pointed out that current large-scale models are not fully controllable in terms of the generated contents.
Figure 16: Changed clothing styles (VITON-HD).
@@ 5.2 Limited Human Parsing
Clothing style is an important aspect to show the performance of virtual try-on, but the current models has poor performance on unconventional styles. From top to bottom in Fig. 15 are short T-shirts, loose-try-on tops, long T-shirts and light tops. Fig. 16 also shows some examples of unconventional styles in VITON-HD dataset. Compared with the correct wearing state of the corresponding clothing in the rightmost column of each row, the try-on results generated by the model have obvious biases, and the model tends to make the clothing fit perfectly on the upper body of the person body, presenting a slimming effect. This is unfair for clothing with loose styles or special designs in length, such as the results in the first and third rows of Fig. 15, where the short T-shirts are forcibly stretched to the normal length of T-shirts, and the long T-shirts are forcibly reduced/truncated to the normal length of T-shirts. Essentially, the model does not really understand the clothing to be tried on, and it only deforms the clothing image and fills it in the original clothing area for replacement. The main reason attributes to that the model is trained with paired data where target clothes and the clothes on person are actually the same one, and in reality it is difficult to collect unpaired data. Clothing-agnostic person representation still deserve exploration.
Figure 17: Artifacts caused by noisy parsing.
@@ 5.3 Limited Clothing Parsing
It is worth mention that clothing parsing is not necessary, e.g., for some large-scale image generation models, where only the character image and clothing image are needed as conditions, and a mask is used to guide the area to be tried on. Due to the strong feature parsing ability and scalability of large image generation models, the dependency on parsing will be also alleviated.
@@ 6 Future Work
Clothing-agnostic person representation. As the training data is paired where the person already wears target clothes, it will harm the generalization ability if person representation contains clothes information. Current representations $\mathcal{P}_{\begin{subarray}{c}3,6,7,10,11\end{subarray}}$ in Fig. 4 contains clues from the original clothes such as clothing contour, which will affect the clothing style of target clothes. Totally deleting the body area (i.e., $\mathcal{P}_{\begin{subarray}{c}4,12\end{subarray}}$ will get rid of original clothes but lose person characteristics such as skin color. Constructing triplets where a person wears different clothes could be a solution, but current methods [47, 31, 40] still rely on parser-based methods to construct the triplet, which inherits the limitation of human parser. Future direction could try to construct training triplets with parser-free approaches such as recent large-scale models, or totally delete original clothes but keep a sampler at exposed skin. For diffusion models based on mask inpainting, how to decouple the mask with original clothes also deserves exploring.
@@
Controllable generation with large-scale model. Large-scale image generation models show powerful generalization and high quality, which is appealing for virtual try-on tasks. However, current models only meet semantic-level control such as satisfying the description from text. More precise and strict control has not been well solved such as preserving the original styles (e.g., Fig. 16). Effectively encoding clothing image and injecting to the generation model, or just using large-scale models to refine the warped clothes deserve a try.
@@
Specialized evaluation metrics. Existing evaluation metrics are not specialized for virtual try-on tasks [8]. In this paper, a new criteria is proposed with advanced pre-trained model and specifically evaluate the semantic information for try-on and non-try-on areas. Future attention could be paid to assessing the clarity and realism of clothing patterns, the preservation of clothing styles and the person characteristics.
@@
Scenario-specific try-on. Real-world try-on is limited to dressing rooms, while try-on effects differ in different scenarios such as indoor or outdoor, different seasons and different occasions. Virtual try-on has advantages in synthesizing the background, which provides imaginary space for users to consider the suitability of target clothes. Furthermore, generating videos to showcase the effects in different scenarios would be more engaging [36].
7 Conclusion
With the rapid development of image generation models, the performance of image-based virtual try-on task is getting closer to practical applications. To the best of our knowledge, this is the first systematic review for image-based virtual try-on with unified evaluation and in-depth analysis. In this review, we first propose taxonomy in aspects of try-on indication, cloth warping and try-on modules and give in-depth analysis towards representative methods in these designs. Afterwards, we introduce datasets and evaluation criteria, and uniformly evaluate representative methods using mainstream and new proposed criteria. Quantitative, qualitative and user study experiments are conducted and analyzed. Finally, we show current unresolved issues and point out future directions. We hope this survey provides an effective way to comprehensively understand image-based virtual try-on and inspires further exploration of this research field.
[…]
Acknowledgements This work was supported in part by the National Natural Science Foundation of China under Grant U21B2024 and Grant 61902277.
Declarations
  • Competing interests. The authors have no competing interests to declare that are relevant to the content of this article.
@@ 7 Conclusion
  • Data availability. The uniformly implemented evaluation metrics, dataset and collected methods will be made public available at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.